# Violation of Luttinger’s theorem in one-dimensional interacting fermions

### v1

@@ I Introduction
Recent studies have shown that the Luttinger theorem is not universally valid [8, 9, 10], especially in systems lacking particle-hole (p-h) symmetry. Notable examples include the Falicov-Kimball model in the strong-coupling limit [8], the generalized $t$-$V$ model [9], the $t$-$J$ model [10], and the Hubbard model [10, 11]. In these systems, strong correlations can cause the Luttinger integral $I_{L}$ to deviate from the actual particle density [11], signaling not only a reconstruction of the Fermi surface but also a fundamental change in the nature of low-energy excitations.
Strongly correlated low-dimensional systems exhibit rich dynamical behavior that goes beyond the Landau paradigm. Key observables such as single-particle spectra, momentum distributions, and correlation functions can be probed via cold-atom experiments [12], X-ray scattering [13], and angle-resolved photoemission spectroscopy (ARPES) [14], serving as powerful tools for investigating unconventional metallic states, metal-insulator transitions [15], and low-energy excitations. Due to its simple structure and rich correlation effects, the 1D generalized $t$-$V$ model provides an ideal platform for studying interaction-driven Fermi surface reconstructions and the breakdown of conventional theories. Although previous works have primarily focused on the half-filled regime [9, 16], a systematic investigation of Green’s functions and momentum distributions away from half-filling, especially in connection with the continuous evolution of correlation mechanisms, is still lacking.
@@
In case (ii), we identify a metallic phase beyond the LL paradigm—a non-Luttinger liquid (NLL) that retains finite spectral weight at $k_{F}$ but lacks the characteristic branch-cut singularity. Zeros of $\operatorname{Re}G(k,0)$ emerge away from $k_{F}$, thereby modifying $I_{L}$ in the absence of topological protection. This breakdown of the LL framework is also reflected in a smooth momentum distribution near $k_{F}$. Identifying and characterizing the NLL phase constitutes a central part of this work and will be discussed in detail below.
@@ II Model and Method
To investigate 1D strongly correlated fermionic systems, LL theory provides an effective low-energy description for gapless metallic phases. However, this framework has known limitations, particularly in capturing critical behavior beyond the LL paradigm in the strong-coupling regime [17, 16]. To explore such regimes, we consider a minimal but versatile model: the generalized $t$-$V$ model (also known as the $t$-$t^{\prime}$-$V$ model). This model exhibits a wide range of low-energy behaviors through tunable parameters, encompassing the free fermion gas, LL phases, and exotic strongly correlated regimes. It offers a well-controlled setting for studying non-FL behavior, highlighting its distinction from conventional FL properties and examining the applicability or potential breakdown of the Luttinger theorem. The Hamiltonian is given by
@@
Figure 1: (a) Schematic illustration of the Hamiltonian for the $t$-$t^{\prime}$-$V$ model. (b) Zero-temperature phase diagram of the $t$-$t^{\prime}$-$V$ model in the $n$-$V$ plane, with the nearest-neighbor hopping set as the energy unit ($t=1$). At low filling or weak interaction, the system resides in a Luttinger liquid (LL) phase. As the density $n$ and interaction $V/t$ increase beyond a critical threshold, the system enters a non-Luttinger liquid (NLL) phase. At half-filling, an interaction-driven metal-insulator transition drives the system into a charge-density-wave (CDW) phase. Beyond half-filling, the system evolves back into the NLL phase.
@@
[…]
@@ III Numerical Results
For the $t$-$t^{\prime}$-$V$ model, at half-filling, the system opens a gap and enters a charge-density-wave (CDW) phase as the interaction strength increases. At other fillings, extrapolation results indicate that the system remains metallic. For example, at a fermion density $n\approx 0.48$, the system remains gapless even at strong coupling $V/t=10$.
@@
It should be emphasized that the phase boundaries shown in Fig. 1 reflect crossover behavior in a finite-size system and do not correspond to sharply defined phase transitions in the thermodynamic limit. Nevertheless, the qualitative LL–NLL–CDW evolution revealed in this diagram agrees well with the scaling analyses discussed below and thus provides important physical insights. We find that under strong coupling at half-filling, the system does exhibit deviations from the Luttinger theorem. However, these deviations are largely confined to the region near half-filling ($|n-0.5|<0.03$), especially on the over-half-filled side, where the electronic behavior becomes significantly more intricate. This suggests that the emergence of NLL features is not governed solely by electron-electron interactions. Further calculations at different values of $t^{\prime}/t$ reveal a power-law relationship between the degree of deviation and the p-h asymmetry of the system, as clearly demonstrated in Fig. 2.
@@
To verify the reliability of our numerical results, we compared DMRG data with exact diagonalization (ED) results for small systems, finding excellent agreement within the comparable parameter regimes (see Appendix). Our findings are consistent with the ED extrapolations reported in Ref. [9] for systems with $N=14$-$30$ sites. The main improvement in this work is the use of larger systems ($N=82,86,90,94,98,102$), which reduce finite-size effects and significantly improve computational accuracy, particularly addressing previous limitations in the intermediate coupling regime $42<V/t<$ 4 where momentum resolution was constrained.
We also computed $G(k,\omega=0)$ at half-filling for $t^{\prime}/t=0.2,0.4$, extrapolating the Luttinger momentum $k_{L}$ [22] to the thermodynamic limit, as shown in Fig. 2. For the half-filled system with $t^{\prime}=0$, due to p-h symmetry, the real part of the Green’s function satisfies the antisymmetry relation
$\mathrm{Re}\,G(k,0)=-\mathrm{Re}\,G(-k,0),$ (8)
which ensures that regions with $\mathrm{Re}\,G(k,0)>0$ and $<0$ are strictly symmetric in momentum space, leading to $I_{L}=n$. This result holds regardless of interaction strength and directly reflects the p-h symmetry of the system [23]. The main source of error in the data points arises from uncertainties in the finite-size extrapolation process. The fitting curves for $t^{\prime}/t=0.2,0.4$ clearly demonstrate a power-law correlation between the deviation of $I_{L}$ and the degree of p-h asymmetry.
@@ IV Deviation Analysis
$C_{NN}(k)=\frac{1}{L}\sum_{j,\ell}e^{ik(j-\ell)}\left(\langle\hat{n}_{j}\hat{n}_{\ell}\rangle-\langle\hat{n}_{j}\rangle\langle\hat{n}_{\ell}\rangle\right).$ (13)
@@
A closer comparison with $\mathrm{Re}\,G(k,0)$ in Fig. 4 reveals that as long as a singular structure is preserved near $k_{F}$, the correlation function $C_{NN}(k)$ typically retains a prominent feature at $2k_{F}$, reflecting the underlying Luttinger liquid-like behavior. However, in the NLL and CDW regimes, the zero-crossing behavior of $\mathrm{Re}\,G(k,0)$ often lacks the necessary singular support or disrupts the global topological structure of $G(k,0)$ in momentum space, leading to a gradual suppression or smearing of the $2k_{F}$ peak.
@@ V Summary and Discussion
A more precise description is that in the weak-coupling limit, LL and FL phases satisfy the Luttinger theorem due to emergent p-h symmetry. In the strong-coupling regime, $G(k,0)$ often develop zeros at certain momentum where $\operatorname{Re}G(k,0)=0$, implying a divergent self-energy $\Sigma(k,0)$. In such cases, the Luttinger-Ward functional, which generates the free energy from $\Sigma$, becomes ill-defined. Nevertheless, this divergence does not necessarily invalidate the theorem. For systems with p-h symmetry, $G(k,0)$ and related quantities are often (anti)symmetric about $k_{F}$, thus preserving the total Fermi volume. Consequently, the presence of poles, branch cuts, or even nonsingular zeros does not affect the validity of the theorem as long as symmetry is maintained. In symmetry-broken systems, however, smooth zero crossings in $\operatorname{Re}G(k,0)$ can still lead to divergent self-energies, and more critically, the absence of topological constraints allows zero points to drift, ultimately invalidating the theorem.
@@ Acknowledgements.
To verify the accuracy of our DMRG simulations, we benchmarked the results against ED for small system sizes where ED is feasible. We considered the $t$-$t^{\prime}$-$V$ model with PBC at various fermion densities. Specifically, we computed the single-particle Green’s function $G(k,\omega=0)$ for $N_{f}/N=11/26$, $V/t=7$, and $t^{\prime}/t=0,0.2,0.4$ using both methods. We employed a DMRG truncation threshold of $10^{-10}$ and added noise terms to avoid getting trapped in local minima. The sweeping procedure was terminated once the energy variation dropped below $10^{-14}$, ensuring convergence.

### v2

@@ I Introduction
Recent studies have shown that the Luttinger theorem is not universally valid [8, 9, 10, 11, 12], particularly in systems lacking particle-hole (p-h) symmetry. Notable examples include the Falicov-Kimball model in the strong-coupling limit [8], the generalized $t$-$V$ model [9], the $t$-$J$ model [10], and the Hubbard model [10, 11]. In these systems, strong correlations can lead to deviations of the Luttinger integral $I_{L}$ from the particle density [11]. This indicates not only a reconstruction of the Fermi surface but also a fundamental change in low-energy excitations. For instance, Ref. [9] demonstrates a breakdown of the Luttinger theorem in the insulating phase of the half-filled generalized $t$-$V$ model across a metal-insulator transition. Ref. [10] identifies a “Luttinger-breaking” regime near half-filling in the two-dimensional Fermi-Hubbard model in the absence of long-range order. However, exact diagonalization (ED) suffers from the exponential growth of the Hilbert space, while determinant quantum Monte Carlo (DQMC) is hindered by the sign problem [13], which restricts both the accessible parameter space and the attainable numerical accuracy.
Strongly correlated low-dimensional systems exhibit rich dynamical behavior that goes beyond the Landau paradigm. Key observables such as single-particle spectra, momentum distributions, and correlation functions can be probed via cold-atom experiments [14], X-ray scattering [15], and angle-resolved photoemission spectroscopy (ARPES) [16]. These techniques serve as powerful tools for investigating unconventional metallic states, metal-insulator transitions [17], and low-energy excitations. Although LL theory has successfully described many 1D metallic systems, it is based on several strong assumptions, particularly the linearization of the dispersion relation and the expansion of two-body interactions for small momentum transfers near $k_{F}$. In realistic strongly correlated systems, these approximations may break down, as symmetry breaking and renormalization of excitation modes due to strong interactions can lead to significant deviations from conventional LL behavior. To date, no rigorous mathematical proof has settled this question. It is thus essential to explore whether 1D interacting fermion systems can host metallic states that retain conventional metallic behavior while violating the key assumptions of LL theory. The 1D generalized $t$-$V$ model, with tunable interactions and adjustable p-h asymmetry, provides a natural platform for this investigation.
@@
In case (ii), we identify a metallic phase beyond the LL paradigm, which we term a non-Luttinger liquid (NLL). It arises from the interplay of interaction-driven reconstruction of excitation modes and breaking of p-h symmetry, leading to a topological reorganization of the Fermi points. In the strongly interacting regime near half filling, a new type of 1D metallic state emerges. Its momentum distribution is smooth around $k_{F}$, and the spectral properties deviate significantly from those of the LL. A finite spectral weight persists at $k_{F}$, while the characteristic branch-cut singularities are gradually destroyed. Zeros of $\operatorname{Re}G(k,0)$ appear away from $k_{F}$, which, in the absence of topological protection, strongly affect $I_{L}$. This phase combines pronounced strong-coupling features with a reconstructed topological structure, revealing a class of non-Luttinger 1D correlated metals that, to our knowledge, has not been systematically explored in the literature. Identifying and characterizing this NLL phase constitutes a central part of the present work and will be discussed in detail below.
@@ II Model and Method
To investigate 1D strongly correlated fermionic systems, LL theory provides an effective low-energy description for gapless metallic phases. However, this framework has known limitations, particularly in capturing critical behavior beyond the LL paradigm in the strong-coupling regime [18, 19]. To explore such regimes, we consider a minimal but versatile model: the generalized $t$-$V$ model (also known as the $tt^{\prime}-$$V$ model). This model exhibits a wide range of low-energy behaviors through tunable parameters, encompassing the free fermion gas, LL phases, and exotic strongly correlated regimes. It offers a well-controlled setting for studying non-FL behavior, highlighting its distinction from conventional FL properties and examining the applicability or potential breakdown of the Luttinger theorem. The Hamiltonian is given by
@@
Figure 1: (a) Schematic illustration of the Hamiltonian for the $tt^{\prime}-$$V$ model. (b) Zero-temperature phase diagram of the $tt^{\prime}-$$V$ model in the $n$-$V$ plane, with the nearest-neighbor hopping set as the energy unit ($t=1$). At low filling or weak interaction, the system resides in a Luttinger liquid (LL) phase. As the density $n$ and interaction $V/t$ increase beyond a critical threshold, the system enters a non-Luttinger liquid (NLL) phase. At half-filling, an interaction-driven metal-insulator transition drives the system into a charge-density-wave (CDW) phase. Beyond half-filling, the system evolves back into the NLL phase. (c)Non-interacting dispersion relation and the corresponding density of states (DOS) at half-filling, under the same parameters as used in (b). The inclusion of $t^{\prime}$ shifts the position of the band maxima in the Brillouin zone and results in a clearly asymmetric DOS with respect to zero energy.
@@
$G(k,\omega)=-i\int_{0}^{\infty}dt\,e^{i(\omega+\mu)t}\langle 0|\{\hat{c}_{k}^{\dagger},\,\hat{c}_{k}(t)\}_{+}|0\rangle,$ (3)
@@ III Numerical Results
For the $tt^{\prime}-$$V$ model, at half-filling, the system opens a gap and enters a charge-density-wave (CDW) phase as the interaction strength increases. At other fillings, extrapolation results indicate that the system remains metallic. For example, at a fermion density $n\approx 0.48$, the system remains gapless even at strong coupling $V/t=10$.
@@
It should be emphasized that the phase boundaries shown in Fig. 1 reflect crossover behavior in a finite-size system and do not correspond to sharply defined phase transitions in the thermodynamic limit. Nevertheless, the qualitative LL–NLL–CDW evolution revealed in this diagram agrees well with the scaling analyses discussed below and thus provides important physical insights. We find that under strong coupling at half-filling, the system does exhibit deviations from the Luttinger theorem. However, these deviations are largely confined to the region near half-filling (|n0.5|<0.03|n-0.5|<0.03), especially on the over-half-filled side, where the electronic behavior becomes significantly more intricate. This suggests that the emergence of NLL features is not governed solely by electron-electron interactions. Further calculations at different values of $t^{\prime}/t$ reveal a power-law relationship between the degree of deviation and the p-h asymmetry of the system, as clearly demonstrated in Fig. 2.
@@
To verify the reliability of our numerical results, we compared DMRG data with ED results for small systems, finding excellent agreement within the comparable parameter regimes (see Appendix). Our findings are consistent with the ED extrapolations reported in Ref. [9] for systems with N=14N=14-$30$ sites. The main improvement in this work is the use of larger systems ($N=82,86,90,94,98,102$), which reduce finite-size effects and significantly improve computational accuracy, particularly addressing previous limitations in the intermediate coupling regime $2<V/t<4$ where momentum resolution was constrained.
We also computed $G(k,\omega=0)$ at half-filling for $t^{\prime}/t=0.2,0.4$, extrapolating the Luttinger momentum $k_{L}$ [24] to the thermodynamic limit, as shown in Fig. 2. For the half-filled system with $t^{\prime}=0$, due to the exact p-h symmetry, the real part of the Green’s function satisfies the relation
$\mathrm{Re}\,G(k-k_{F},0)=-\mathrm{Re}\,G(k_{F}-k,0),$ (8)
which ensures that regions with $\mathrm{Re}\,G(k,0)>0$ and <0<0 are strictly antisymmetric in momentum space about $k_{F}$, leading to $I_{L}=n$. This result holds regardless of interaction strength and directly reflects the p-h symmetry of the model [25]. When a next-nearest-neighbor hopping $t^{\prime}\neq 0$ is introduced, the exact p-h symmetry is broken, and Eq. (8) no longer holds strictly. Nevertheless, in the low-energy LL regime, emergent p-h and Lorentz symmetries appear, so that Eq. (8) remains approximately valid. The main source of uncertainty in the data points comes from the finite-size extrapolation process. The fitting curves for $t^{\prime}/t=0.2,0.4$ clearly show a power-law correlation between the deviation of $I_{L}$ and the degree of p-h asymmetry.
@@ IV Deviation Analysis
[…]
@@
A closer comparison with $\mathrm{Re}\,G(k,0)$ in Fig. 4 reveals that as long as a singular structure is preserved near $k_{F}$, the correlation function $C_{NN}(k)$ typically retains a prominent feature at $2k_{F}$, reflecting the underlying Luttinger liquid-like behavior. However, in the NLL and CDW regimes, the zero-crossing behavior of $\mathrm{Re}\,G(k,0)$ often lacks the necessary singular support or disrupts the global topological structure of $G(k,0)$  [28, 29] in momentum space, leading to a gradual suppression or smearing of the $2k_{F}$ peak.
@@ V Summary and Discussion
A more precise description is that in the weak-coupling limit, LL and FL phases satisfy the Luttinger theorem due to emergent p-h symmetry. In the strong-coupling regime, $G(k,0)$ often develop zeros at certain momenta where $\operatorname{Re}G(k,0)=0$, implying a divergent self-energy $\Sigma(k,0)$. In such cases, the Luttinger-Ward functional, which generates the free energy from $\Sigma$, becomes ill-defined. Nevertheless, this divergence does not necessarily invalidate the theorem. For systems with p-h symmetry, $G(k,0)$ and related quantities are often (anti)symmetric about $k_{F}$, thus preserving the total Fermi volume. Consequently, the presence of poles, branch cuts, or even nonsingular zeros does not affect the validity of the theorem as long as symmetry is maintained. In symmetry-broken systems, however, smooth zero crossings in $\operatorname{Re}G(k,0)$ can still lead to divergent self-energies, and more critically, the absence of topological constraints allows zero points to drift, ultimately invalidating the theorem. We propose that the interplay between the evolution of excitation modes and symmetry breaking breaks the original topological constraints, leading to a shift in the position of the generalized Fermi surface and its topological reconstruction, thereby driving the system away from the LL fixed point toward a new phase. This cooperative effect is not merely a simple superposition of the two mechanisms, but could represent a genuinely new type of collective phenomenon whose physical implications have not yet been systematically explored.
@@ Acknowledgements.
The DMRG method used in this work is a high-precision algorithm for ground-state simulations. The wavefunction is represented in the MPS form, and local optimizations combined with singular value decomposition (SVD) are applied to truncate the reduced density matrix, retaining only the most entangled states [20]. This enables an efficient compression of the Hilbert space. While ED suffers from exponential scaling with system size, DMRG exploits the area law of entanglement entropy in 1D systems, reducing the computational cost to polynomial order (typically $O(\chi^{3})$, with $\chi$ the MPS bond dimension). In practice, we systematically increase $\chi$ until the ground-state energy and relevant observables converge within the desired tolerance.
To verify the accuracy of our DMRG simulations, we benchmarked them against ED on small systems where the latter is feasible. Specifically, we considered the $tt^{\prime}-$$V$ model with PBC at various fermion densities, and computed the single-particle Green’s function $G(k,0)$ for $N_{f}/N=11/26$, $V/t=7$, and $t^{\prime}/t=0,0.2,0.4$. In the DMRG simulations, a truncation threshold of $10^{-10}$ was imposed, and small noise terms were introduced to avoid trapping in local minima. The sweeps were terminated once the energy change fell below $10^{-13}$, ensuring global convergence of the wavefunction and the physical reliability of the results.